# Learning to Explore with Lagrangians for Bandits under Unknown Linear Constraints

### v1

@@ G.4 Upper Bound for LAGEX
The expected sample complexity of LAGEX satisfies
$\mathbb{E}[\tau]\leq T_{0}(\delta)+2\mathfrak{s}+C{K}$
where $T_{0}(\delta)\mathrel{\triangleq}\max\left\{t\in\mathbb{N}:t\leq T_{\mathcal{F}}({\boldsymbol{\mu}})\left(\beta(t,\delta)+\mathcal{O}\left(\sqrt{t\log t}\right)+\mathcal{O}\left(\sqrt{Qt}\right)\right)\right\}$ and $C$ is a problem independent constant.
@@
$G_{t}=\{\forall t\leq T,\forall a\in[{K}]:N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\mu}}_{a})\leq g(t)\wedge\|(\tilde{\mathbf{A}}_{t}-\mathbf{A})\boldsymbol{\omega}\|_{\infty}\leq\rho(t,\delta)\}$
@@
$\sum_{s=1}^{T}\mathbb{P}\left(G_{t}^{c}\right)$ $=\sum_{s=1}^{T}\mathbb{P}\left(\sum_{a=1}^{K}N_{a,s}d(\hat{{\boldsymbol{\mu}}}_{a,s},{\boldsymbol{\mu}}_{a})>g(s)\right)+\sum_{s=1}^{T}\mathbb{P}\left(\|(\tilde{\mathbf{A}}_{s}-\mathbf{A})\boldsymbol{\omega}\|_{\infty}>\rho(s,\delta)\forall\boldsymbol{\omega}\in\Delta_{K}\right)$
@@
$\sum_{s=1}^{T}\mathbb{P}\left(\sum_{a=1}^{K}N_{a,s}d(\hat{{\boldsymbol{\mu}}}_{a,s},{\boldsymbol{\mu}}_{a})>g(s)\right)\leq 1+\sum_{s=2}^{\infty}\frac{\exp(2)}{s^{3}\log s}(g(s)+g^{2}(s)\log s)\leq 19.48$
as a direct consequence of (Degenne et al., 2019a, , Lemma 6). The second cumulative probability is bounded by $\zeta(1)<0.578$ using Lemma 4, which is finite.
In the next step, we work with the stopping criterion where we do not have access to $\mathcal{F}$ rather a bigger feasible set ^$\hat{\mathcal{F}}_{t}$.
@@
$\beta(t,\delta)\geq\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t},\boldsymbol{\pi})}\sum_{a=1}^{K}N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})\,,$
@@
$\beta(t,\delta)\geq\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t},\boldsymbol{\pi})}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})-(1+\sqrt{t}){K}$ (19)
@@
$\beta(t,\delta)\geq\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t},\boldsymbol{\pi})}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})-(1+\sqrt{t}){K}-\mathcal{O}(\sqrt{t\log t})$ (21)
@@
$\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t},\boldsymbol{\pi})}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{s,a}d(\hat{{\boldsymbol{\mu}}}_{a,s},{\boldsymbol{\lambda}}_{a})$
$\geq$ $\max_{\boldsymbol{\omega}\in\hat{\mathcal{F}}_{t}}\sum_{a=1}^{K}\sum_{s=1}^{t}\boldsymbol{\omega}_{a}U_{a,s}-R_{t}^{\boldsymbol{\omega}}-\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}C_{a,s}+\boldsymbol{l}_{t}^{\top}\tilde{\mathbf{A}}_{t}\boldsymbol{\omega}_{t}\,,$ (22)
@@
$R_{{\boldsymbol{\lambda}}}^{t}\mathrel{\triangleq}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}d(\hat{{\boldsymbol{\mu}}}_{a,s},\lambda_{a,s})-\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t},\boldsymbol{\pi})}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}d(\hat{{\boldsymbol{\mu}}}_{a,s},{\boldsymbol{\lambda}}_{a})\leq 0$
@@
$\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t},\boldsymbol{\pi})}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}d(\hat{{\boldsymbol{\mu}}}_{a,s},{\boldsymbol{\lambda}}_{a})\geq\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}U_{a,s}-\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}C_{a,s}\,.$
@@
$\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}},\boldsymbol{\pi})}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}d(\hat{{\boldsymbol{\mu}}}_{a,s},\lambda_{s})\geq\max_{\boldsymbol{\omega}\in\hat{\mathcal{F}}}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a}U_{a,s}-R_{t}^{\boldsymbol{\omega}}-\underset{T1}{\underbrace{\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}C_{a,s}}}+\boldsymbol{l}_{t}^{\top}\tilde{\mathbf{A}}_{t}\boldsymbol{\omega}_{t}$
$\geq\max_{\boldsymbol{\omega}\in\hat{\mathcal{F}}}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a}U_{a,s}-\mathcal{O}(\sqrt{Qt})-\underset{T1}{\underbrace{\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}C_{a,s}}}+\boldsymbol{l}_{t}^{\top}\tilde{\mathbf{A}}_{t}\boldsymbol{\omega}_{t}$
which shows that LAGEX is a approximate saddle point algorithm with a slack $R_{t}^{\boldsymbol{\omega}}+\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}C_{a,s}-\boldsymbol{l}_{t}\tilde{\mathbf{A}}_{t}\boldsymbol{\omega}_{t}$.
@@
$\beta(t,\delta)$ $\geq\max_{\boldsymbol{\omega}\in\Delta_{K}}\left(\sum_{a=1}^{K}\sum_{s=1}^{t}\boldsymbol{\omega}_{a}d({\boldsymbol{\mu}}_{a},\lambda_{a,s})\right)-\mathcal{O}(\sqrt{Qt})-\mathcal{O}(\sqrt{t\log t})-(1+\sqrt{t}){K}+\boldsymbol{l}_{t}^{\top}\tilde{\mathbf{A}}_{t}\boldsymbol{\omega}_{t}$
$\geq\max_{\boldsymbol{\omega}\in\Delta_{K}}\sum_{a=1}^{K}\sum_{s=1}^{t}\boldsymbol{\omega}_{a}d({\boldsymbol{\mu}}_{a},\lambda_{a,s})-\mathcal{O}(\sqrt{Qt})-\mathcal{O}(\sqrt{t\log t})-(1+\sqrt{t}){K}+\boldsymbol{l}_{t}^{\top}\tilde{\mathbf{A}}_{t}\boldsymbol{\omega}_{t}$
@@
The last inequality is true due to Lemma 6, i.e. $-\boldsymbol{l}_{t}^{\top}\tilde{\mathbf{A}}_{t}\boldsymbol{\omega}_{t}\leq\frac{\psi_{t}}{T_{\hat{\mathcal{F}}_{t}}({\boldsymbol{\mu}})}$.
@@
$\max_{\boldsymbol{\omega}\in\Delta_{K}}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a}d({\boldsymbol{\mu}}_{a},\lambda_{a,s})\geq t\max_{\boldsymbol{\omega}\in\Delta_{K}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t},\boldsymbol{\pi})}\sum_{a=1}^{K}\boldsymbol{\omega}_{a}d({\boldsymbol{\mu}}_{a},{\boldsymbol{\lambda}}_{a})=tT_{\hat{\mathcal{F}}_{t}}^{-1}({\boldsymbol{\mu}})$
@@
since $\mathrm{ch}(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\mu}}_{a})\leq d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\mu}}_{a})$. At time $s\in\mathbb{N}$, let $\boldsymbol{\pi}^{\prime}$ be an extreme point in ^$\hat{\mathcal{F}}_{s}$ that is not the r-good optimal policy. But since it is an extreme point in ^$\hat{\mathcal{F}}_{s}$ that shares $({K}-1)$ active constraints with $\boldsymbol{\pi}^{\star}$, it has to be an r-good optimal policy w.r.t some ${\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{s}}(\hat{{\boldsymbol{\mu}}}_{s},\boldsymbol{\pi}):\boldsymbol{\pi}^{\prime}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{s}}^{r}}{\boldsymbol{\lambda}}^{\top}\boldsymbol{\pi}$. So we again define the event $B_{t}$ as
$B_{t}\mathrel{\triangleq}\bigg{\{}{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t},\boldsymbol{\pi}):\boldsymbol{\pi}^{\prime}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}{\boldsymbol{\lambda}}^{\top}\boldsymbol{\pi}\notin\Pi_{\mathcal{F}}^{r}\bigg{\}}$
@@
Now to break the RHS of the above inequality we go back to step 5 of the proof of part (a) where we showed LAGEX is an approximate saddle point algorithm. In this case the slack will be $x_{t}=R_{n_{\boldsymbol{\pi}^{\prime}}(t)}^{\boldsymbol{\omega}}+\sum_{s=1,\boldsymbol{\pi}_{s}=\boldsymbol{\pi}^{\prime}}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}C_{a,s}-\boldsymbol{l}_{t}\tilde{\mathbf{A}}_{t}\boldsymbol{\omega}_{t}$. Therefore we can write the RHS of Equation (28) as
@@
$\geq$ $\max_{\boldsymbol{\pi}\in\hat{\mathcal{F}}_{s}}\underbrace{\sum_{s=1,\boldsymbol{\pi}_{s}=\boldsymbol{\pi}^{\prime}}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}U_{a,s}-R_{n_{\boldsymbol{\pi}^{\prime}}(t)}^{\boldsymbol{\omega}}}_{T1}-\underbrace{\sum_{s=1,\boldsymbol{\pi}_{s}=\boldsymbol{\pi}^{\prime}}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}C_{a,s}}_{T2}+\boldsymbol{l}_{t}^{\top}\tilde{\mathbf{A}}_{t}\boldsymbol{\omega}_{t}\,.$ (29)
@@
as ^$\hat{\mathcal{F}}_{t}\to\mathcal{F}$ and $\psi_{t}\to\mathfrak{s}$, we have the explicit upper bound
$\mathbb{E}[\tau]\leq T_{0}(\delta)+2\mathfrak{s}+C{K}$
where $T_{0}(\delta)\mathrel{\triangleq}\max\left\{t\in\mathbb{N}:t\leq T_{\mathcal{F}}({\boldsymbol{\mu}})\left(\beta(t,\delta)+\mathcal{O}\left(\sqrt{t\log t}\right)+\mathcal{O}\left(\sqrt{Qt}\right)\right)\right\}$ and $C$ is a problem independent constant that appears due to the concentration guaranty to good event $G_{t}$ i.e concentration of mean vector and constraint matrix.

### v2

@@ G.4 Upper Bound for LAGEX
The expected sample complexity of LAGEX $\lim_{\begin{subarray}{c}\delta\rightarrow 0\end{subarray}}\frac{\mathbb{E}[\tau_{\delta}]}{\log(1/\delta)}\leq T_{\mathcal{F},r}({\boldsymbol{\mu}}).$
@@
$G_{t}=\{\forall t\leq T,\forall a\in[{K}]:N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\mu}}_{a})\leq g(t)\wedge\|(\tilde{A}_{t}-A)\boldsymbol{\omega}\|_{\infty}\leq\rho(t,\delta)\}$
@@
$\sum_{s=1}^{\top}\mathbb{P}\left(G_{t}^{c}\right)$ $=\sum_{s=1}^{\top}\mathbb{P}\left(\sum_{a=1}^{K}N_{a,s}d(\hat{{\boldsymbol{\mu}}}_{a,s},{\boldsymbol{\mu}}_{a})>g(s)\right)+\sum_{s=1}^{\top}\mathbb{P}\left(\|(\tilde{A}_{s}-A)\boldsymbol{\omega}\|_{\infty}>\rho(s,\delta)\forall\boldsymbol{\omega}\in\Delta_{K}\right)$
@@
$\sum_{s=1}^{\top}\mathbb{P}\left(\sum_{a=1}^{K}N_{a,s}d(\hat{{\boldsymbol{\mu}}}_{a,s},{\boldsymbol{\mu}}_{a})>g(s)\right)\leq 1+\sum_{s=2}^{\infty}\frac{\exp(2)}{s^{3}\log s}(g(s)+g^{2}(s)\log s)\leq 19.48$
as a direct consequence of (Lemma 6 of Degenne et al., 2019a ). The second cumulative probability is bounded by $\zeta(1)<0.578$ using Lemma 4, which is finite.
In the next step, we work with the stopping criterion where we do not have access to $\mathcal{F}$ rather a bigger feasible set $\hat{\mathcal{F}}_{t}$.
@@
$\beta(t,\delta)\geq\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})}\sum_{a=1}^{K}N_{a,t}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})\,$,
@@
$\beta(t,\delta)\geq\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})-(1+\sqrt{t}){K}$ (17)
@@
$\beta(t,\delta)\geq\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\lambda}}_{a})-(1+\sqrt{t}){K}-\mathcal{O}(\sqrt{t\log t})$ (19)
@@
$\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{s,a}d(\hat{{\boldsymbol{\mu}}}_{a,s},{\boldsymbol{\lambda}}_{a})$
$\geq$ $\max_{\boldsymbol{\omega}\in\hat{\mathcal{F}}_{t}}\sum_{a=1}^{K}\sum_{s=1}^{t}\boldsymbol{\omega}_{a}U_{a,s}-R_{t}^{\boldsymbol{\omega}}-\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}C_{a,s}+\boldsymbol{l}_{t}^{\top}\tilde{A}_{t}\boldsymbol{\omega}_{t}\,$, (20)
@@
$R_{{\boldsymbol{\lambda}}}^{t}\mathrel{\triangleq}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}d(\hat{{\boldsymbol{\mu}}}_{a,s},\lambda_{a,s})-\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}d(\hat{{\boldsymbol{\mu}}}_{a,s},{\boldsymbol{\lambda}}_{a})\leq 0$
@@
$\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}d(\hat{{\boldsymbol{\mu}}}_{a,s},{\boldsymbol{\lambda}}_{a})\geq\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}U_{a,s}-\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}C_{a,s}\,.$
@@
$\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}}({\boldsymbol{\mu}})}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}d(\hat{{\boldsymbol{\mu}}}_{a,s},\lambda_{s})\geq\max_{\boldsymbol{\omega}\in\hat{\mathcal{F}}}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a}U_{a,s}-R_{t}^{\boldsymbol{\omega}}-\underset{T1}{\underbrace{\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}C_{a,s}}}+\boldsymbol{l}_{t}^{\top}\tilde{A}_{t}\boldsymbol{\omega}_{t}$
$\geq\max_{\boldsymbol{\omega}\in\hat{\mathcal{F}}}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a}U_{a,s}-\mathcal{O}(\sqrt{Qt})-\underset{T1}{\underbrace{\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}C_{a,s}}}+\boldsymbol{l}_{t}^{\top}\tilde{A}_{t}\boldsymbol{\omega}_{t}$
which shows that LAGEX is a approximate saddle point algorithm with a slack $R_{t}^{\boldsymbol{\omega}}+\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}C_{a,s}-\boldsymbol{l}_{t}\tilde{A}_{t}\boldsymbol{\omega}_{t}$.
@@
$\beta(t,\delta)$ $\geq\max_{\boldsymbol{\omega}\in\Delta_{K}}\left(\sum_{a=1}^{K}\sum_{s=1}^{t}\boldsymbol{\omega}_{a}d({\boldsymbol{\mu}}_{a},\lambda_{a,s})\right)-\mathcal{O}(\sqrt{Qt})-\mathcal{O}(\sqrt{t\log t})-(1+\sqrt{t}){K}+\boldsymbol{l}_{t}^{\top}\tilde{A}_{t}\boldsymbol{\omega}_{t}$
$\geq\max_{\boldsymbol{\omega}\in\Delta_{K}}\sum_{a=1}^{K}\sum_{s=1}^{t}\boldsymbol{\omega}_{a}d({\boldsymbol{\mu}}_{a},\lambda_{a,s})-\mathcal{O}(\sqrt{Qt})-\mathcal{O}(\sqrt{t\log t})-(1+\sqrt{t}){K}+\boldsymbol{l}_{t}^{\top}\tilde{A}_{t}\boldsymbol{\omega}_{t}$
@@
The last inequality is true due to Lemma 6, i.e. $\boldsymbol{l}_{t}^{\top}\tilde{A}_{t}\boldsymbol{\omega}_{t}\leq\frac{\psi_{t}}{T_{\hat{\mathcal{F}}_{t}}({\boldsymbol{\mu}})}$.
@@
$\max_{\boldsymbol{\omega}\in\Delta_{K}}\sum_{s=1}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a}d({\boldsymbol{\mu}}_{a},\lambda_{a,s})\geq t\max_{\boldsymbol{\omega}\in\Delta_{K}}\max_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}\inf_{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t})}\sum_{a=1}^{K}\boldsymbol{\omega}_{a}d({\boldsymbol{\mu}}_{a},{\boldsymbol{\lambda}}_{a})=tT_{\hat{\mathcal{F}}_{t}}^{-1}({\boldsymbol{\mu}})$
@@
since $\mathrm{ch}(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\mu}}_{a})\leq d(\hat{{\boldsymbol{\mu}}}_{a,t},{\boldsymbol{\mu}}_{a})$. At time $s\in\mathbb{N}$, let $\boldsymbol{\pi}^{\prime}$ be an extreme point in $\hat{\mathcal{F}}_{s}$ that is not the r-good optimal policy. But since it is an extreme point in $\hat{\mathcal{F}}_{s}$ that shares (K1)({K}-1) active constraints with $\boldsymbol{\pi}^{\star}$, it has to be an r-good optimal policy w.r.t some ${\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{s}}(\hat{{\boldsymbol{\mu}}}_{s},\boldsymbol{\pi}):\boldsymbol{\pi}^{\prime}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{s}}^{r}}{\boldsymbol{\lambda}}^{\top}\boldsymbol{\pi}$. So we again define the event $B_{t}$ as
$B_{t}\mathrel{\triangleq}\bigg\{{\boldsymbol{\lambda}}\in\Lambda_{\hat{\mathcal{F}}_{t}}(\hat{{\boldsymbol{\mu}}}_{t}):\boldsymbol{\pi}^{\prime}=\operatorname*{arg\,max}_{\boldsymbol{\pi}\in\Pi_{\hat{\mathcal{F}}_{t}}^{r}}{\boldsymbol{\lambda}}^{\top}\boldsymbol{\pi}\notin\Pi_{\mathcal{F}}^{r}\bigg\}$
@@
Now to break the RHS of the above inequality we go back to step 5 of the proof of part (a) where we showed LAGEX is an approximate saddle point algorithm. In this case the slack will be $x_{t}=R_{n_{\boldsymbol{\pi}^{\prime}}(t)}^{\boldsymbol{\omega}}+\sum_{s=1,\boldsymbol{\pi}_{s}=\boldsymbol{\pi}^{\prime}}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}C_{a,s}-\boldsymbol{l}_{t}\tilde{A}_{t}\boldsymbol{\omega}_{t}$. Therefore we can write the RHS of Equation (26) as
@@
$\geq$ $\max_{\boldsymbol{\pi}\in\hat{\mathcal{F}}_{s}}\underbrace{\sum_{s=1,\boldsymbol{\pi}_{s}=\boldsymbol{\pi}^{\prime}}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}U_{a,s}-R_{n_{\boldsymbol{\pi}^{\prime}}(t)}^{\boldsymbol{\omega}}}_{T1}-\underbrace{\sum_{s=1,\boldsymbol{\pi}_{s}=\boldsymbol{\pi}^{\prime}}^{t}\sum_{a=1}^{K}\boldsymbol{\omega}_{a,s}C_{a,s}}_{T2}+\boldsymbol{l}_{t}^{\top}\tilde{A}_{t}\boldsymbol{\omega}_{t}\,.$ (27)
@@
as $\hat{\mathcal{F}}_{t}\to\mathcal{F}$ and $\psi_{t}\to\mathfrak{s}$, we have the explicit upper bound
$\mathbb{E}[\tau]\leq T_{0}(\delta)+2\mathfrak{s}+C{K}$ (28)
where $T_{0}(\delta)\mathrel{\triangleq}\max\left\{t\in\mathbb{N}:t\leq T_{\mathcal{F}}({\boldsymbol{\mu}})\left(\beta(t,\delta)+\mathcal{O}\left(\sqrt{t\log t}\right)+\mathcal{O}\left(\sqrt{Qt}\right)\right)\right\}$ and $C$ is a problem independent constant that appears due to the concentration guaranty to good event $G_{t}$ i.e concentration of mean vector and constraint matrix. Thus, dividing both sides by $\log(\frac{1}{\delta})$ and taking the limit $\delta\rightarrow 0$ we conclude the proof. Note that, we do not see the effect of not knowing the constraints in the asymptotic guaranty, though in non-asymptotic sense in Equation (28), we see an additive effect of the shadow price on the sample complexity. ∎